The following table describes the production possibilities of two cities in the country of Basebalia:

 

 

Wheat per Worker per Hour Corn per Worker per Hour

Boston 6 6


Chicago 2 4


 

Which city has an absolute advantage in the production of wheat and corn? Which city has a comparative advantage in the production of wheat and corn? Explain your answer.                          


1
Expert's answer
2021-06-11T11:51:46-0400

(i) A city is to have an absolute advantage over another city if it can produce the same good with fewer resources than the other. From the table above; Boston produces 1unit of Wheat/worker in "\frac{1}{6}" worker-hours (10mins) while Chicago requires "\frac{1}{2}" worker-hours (30mins) to produce 1unit of the same (wheat). In corn production, Boston requires "\frac{1}{6}" worker-hours (10mins) to produce 1 unit while Chicago requires "\frac{1}{4}" worker-hours (30mins) to produce a unit of corn. Therefore, Boston city has an absolute advantage over Chicago city.


(ii) A city has a comparative advantage in the production of a good over another city if it has a lower opportunity cost in its production relative to the other city. From the case above; Boston has comparative advantage in wheat production; it gives up 1 unit of corn as opposed to 2 units of corn in Chicago. In corn production; Chicago has comparative advantage since it gives up "\frac{1}{2}" units as opposed 1unit of wheat for Boston.